# HAL: Inducing Human-likeness in LLMs with Alignment

### v1

@@ Abstract
Conversational human-likeness plays a central role in human–AI interaction, yet it has remained difficult to define, measure, and optimize. As a result, improvements in human-like behavior are largely driven by scale or broad supervised training, rather than targeted alignment. We introduce Human Aligning LLMs (HAL), a framework for aligning language models to conversational human-likeness using an interpretable, data-driven reward. HAL derives explicit conversational traits from contrastive dialogue data, combines them into a compact scalar score, and uses this score as a transparent reward signal for alignment with standard preference optimization methods. Using this approach, we align models of varying sizes without affecting their overall performance. In large-scale human evaluations, models aligned with HAL are more frequently perceived as human-like in conversation. Because HAL operates over explicit, interpretable traits, it enables inspection of alignment behavior and diagnosis of unintended effects. More broadly, HAL demonstrates how soft, qualitative properties of language–previously outside the scope for alignment–can be made measurable and aligned in an interpretable and explainable way.
@@ 1 Introduction
Human-like conversational behavior is especially important in settings where interaction quality matters more than task completion. These include role-play and character simulation roleplay1; roleplay2; roleplay3; roleplay4, communication training comm1; comm2; sapien, patient simulation in healthcare and mental health contexts health1; sophie1.0; health2; ptsd; health3, and others.
@@
Recent large-scale Turing test results highlight both progress and limitations llm-turing. GPT-4.5 was judged to be human in 73% of comparisons, while LLaMA-3.1-405B achieved near-chance performance and smaller baselines performed far worse. These findings suggest that scale can improve perceived human-likeness, but they do not explain why, nor do they provide a clear recipe for training models to be more human-like.
In this paper, we introduce Human Aligning LLMs (HAL), a framework for quantifying conversational human-likeness and using it as a reward for alignment. Our approach is entirely data-driven: we extract recurring human-likeness cues from contrastive dialogue data (e.g., Turing tests), compress them into a compact and interpretable set of traits, and combine them into a single scalar score. We then use this score as a reward signal for alignment with standard preference optimization methods such as Direct Preference Optimization (DPO) dpo. Across models of varying sizes, we show that alignment with HAL leads to clear improvements in perceived human-likeness under human evaluation, while largely preserving performance on other benchmarks.
@@ 2.1 Characteristics of Human-likeness From Turing Tests
The Turing test is an imperfect proxy for human-likeness, but it provides two ingredients that are difficult to obtain otherwise: paired dialogues that are designed to be compared, and which side is human. We analyze the Turing test dialogue transcripts released by llm-turing.
@@
Running the LLM-judge over the dataset yields $2735$ total natural-language reason statements, many of which are redundant. We embed each reason statement using a sentence encoder wang2020minilm and cluster them using a density-based clustering algorithm HDBSCAN. We find 53 representative clusters and extract their centroid. The 53 centroid statements still contained redundancies. Hence, we further instruct an LLM to summarize the 53 clusters into distinct Likert-style statements (prompt Appendix Figure 7). This yields a final inventory of 32 characteristics, presented in Appendix Table 5. Henceforth, we refer to these 32 traits as Human-Like 32 Questions or HL32Q.
@@ 3.1 Human-likeness Classifier
Given a dialogue, an LLM judge (HL32Q Judge) observes only the witness responses and rates agreement with each of the 32 statements on a 1–5 Likert scale (prompt Appendix Figure 10). Unlike the Turing test setting, this is not a pairwise comparison. Each dialogue is scored independently, producing a fixed-length feature vector $\mathbf{A}\in\mathbb{R}^{32}$.
@@ 3.2 Feature Reduction and Single Score of Human-likeness
Table 2: Finding characteristics aim to identify the differences between human–human and human–AI data and generate plausible reasons for these differences. The HL32Q judge aims to determine optimal weights for calculating a numerical human-likeness score. Accuracy on the filtered Turing test dataset from llm-turing serves as a proxy for both tasks.
@@ 3.3 Evaluating on OOD Data
Finally, we test whether the HL16Q score generalizes beyond the Turing test data used to derive it. We evaluate on an out-of-distribution dataset consisting of 73 human–human and 73 human–AI dialogues from a separate cancer communication study sophie1.0. This dataset differs in topic, style, and collection procedure (Data proxy in Appendix Table 6).
@@ 4 Inducing Human-likeness with Alignment
We frame this as a preference learning problem. For a given conversational prompt, we generate multiple candidate dialogues, score them with the HL16Q Judge, and construct ranked pairs where the more human-like dialogue is preferred. We use these pairs to align models with Direct Preference Optimization (DPO) dpo.
@@ 4.1 Persona Synthesis
To create diverse yet controlled conversational settings, we synthesize personas that serve as prompts for dialogue generation. We begin with 500 seed personas from the SynthLabs PERSONA dataset synthlabs. These seeds are split into 450 training personas, 25 test personas, and 25 personas reserved for human evaluation. All augmentation is performed after this split to avoid data leakage.
@@ 4.3 Training
We fine-tune seven open-access models ranging from 1B to 72B parameters using DPO dpo. The models were simple instruction-tuned models, with no Mixture-of-Experts (MoE) and Chain-of-Thought (CoT) training, as these add more complexity in alignment training. Unlike standard alignment pipelines, we do not include an intermediate supervised fine-tuning (SFT) step, which typically helps the model learn the data format. All models reliably followed the required dialogue format with prompting alone, making SFT not essential.
We train all models for 10 epochs with DPO using $\beta=0.1$, AdamW optimization in 8-bit precision, a learning rate of $5\times 10^{-5}$, linear scheduling, and a 10% warmup. Training uses an effective batch size of 32 via gradient accumulation, a maximum sequence length of 1024, and no weight decay. We apply LoRA lora with rank 16, $\alpha=32$, and dropout 0.1. All models are trained with 4-bit quantization using HuggingFace Accelerate accelerate in data-parallel mode on 4 NVIDIA H100 GPUs.
@@ 4.4 Training Validation
Despite these improvements, mean HL16Q scores remain negative for most models after training. We attribute this to a domain mismatch between the alignment data and the original Turing test data, for which the bias $b$ was set. We observe the same trait in OOD data (Figure 2) and training data (Appendix Table 8).
@@ 5.1 Chatbot Arena
We adopt a Chatbot Arena–style chatbot-arena interface for pairwise comparison. Using the held-out set of 100 personas, participants interact with two chatbots displayed side by side in random order (Figure 5). For each trial, two models are randomly selected and both receive the same persona prompt. Participants are required to converse with each chatbot for at least two turns before making a decision from 5 choices: Certainly A, Likely A, Tie, Likely B, Certainly B.
@@ Results.
We report both win-rate and Elo scores, using a modified Elo system adapted from Chatbot Arena elo that supports partial wins. Because Elo is sensitive to comparison order, we report the mean Elo score over 500 random shuffles. Full details of the scoring procedure are provided in Appendix A.
@@ 5.2 Emotional Intelligence Benchmarks
To examine whether alignment for human-likeness degrades performance on other capabilities, we evaluate models on two widely used emotional intelligence benchmarks: EmoBench emobench and EQBench3 eqbench; eqbench3_repo_2025. We report results before and after alignment for all models, with full benchmark breakdowns provided in Appendix Table 9.
@@ 7 Limitations
A second limitation concerns the cost of computing HL16Q scores, which requires direct calls to a judge model. Although this cost is incurred only once during data construction, it is significantly higher than training a lightweight reward model, as in RLHF rlhf, or using programmatic rewards in domains such as coding or mathematics grpo. While HAL enables alignment with richer and more nuanced rewards, this judging cost limits scalability at very large scale. In practice, the trade-off between reward expressiveness and computational cost must be considered when deciding whether HAL is suitable for a given alignment task. A promising direction for future work is to distill the HL16Q judge into smaller or ensemble reward models that retain performance while substantially reducing inference cost.
@@ Appendix B Additional Tables
Table 6: OOD dataset Data Proxy, showing structural comparison of Human–AI and Human–Human conversations in a medical setting. Here, the doctor serves as the investigator (I), who interacts with human patient actors and an AI patient as the witness (W). Full data cannot be released due to IRB protection. Although the goal in this study was not to differentiate human witnesses from AI, the pairwise data makes it suitable for our validation. More details about this data at sophie1.0.

### v2

@@ Abstract
Conversational human-likeness plays a central role in human-AI interaction, yet it has remained difficult to define, measure, and optimize. As a result, improvements in human-like behavior are largely driven by scale or broad supervised training, rather than targeted alignment. We introduce Human Aligning LLMs (HAL), a framework for aligning language models to conversational human-likeness using an interpretable, data-driven reward. HAL derives explicit conversational traits from contrastive dialogue data, combines them into a compact scalar score, and uses this score as a transparent reward signal for alignment with standard preference optimization methods. Using this approach, we align models of varying sizes without affecting their overall performance. In large-scale human evaluations, a model aligned with HAL is more frequently perceived as human-like in conversation. Because HAL operates over explicit, interpretable traits, it enables inspection of alignment behavior and diagnosis of unintended effects. More broadly, HAL demonstrates how soft, qualitative properties of language–previously outside the scope for alignment–can be made measurable and aligned in an interpretable and explainable way.
@@ 1 Introduction
Human-like conversational behavior is especially important in settings where interaction quality matters more than task completion. These include role-play and character simulation Shanahan et al. (2023); Wang et al. (2024); Tao et al. (2024); Chen et al. (2024), communication training Yang et al. (2024); Burgues et al. (2024); Hasan et al. (2023), patient simulation in healthcare and mental health contexts Louie et al. (2024); Haut et al. (2025); Elhilali et al. (2025); Baseman et al. (2025); Scherr et al. (2023), and others.
@@
Recent large-scale Turing test results highlight both progress and limitations Jones and Bergen (2025). GPT-4.5 was judged to be human in 73% of comparisons, while LLaMA-3.1-405B achieved near-chance performance and smaller baselines performed far worse. These findings suggest that scale can improve perceived human-likeness, but they do not explain why, nor do they provide a clear recipe for training models to be more human-like.
In this paper, we introduce Human Aligning LLMs (HAL), a framework for quantifying conversational human-likeness and using it as a reward for alignment. Our approach is entirely data-driven: we extract recurring human-likeness cues from contrastive dialogue data (e.g., Turing tests), compress them into a compact and interpretable set of traits, and combine them into a single scalar score. We then use this score as a reward signal for alignment with standard preference optimization methods such as Direct Preference Optimization (DPO) Rafailov et al. (2023). Across models of varying sizes, we show that alignment with HAL leads to clear improvements in perceived human-likeness under human evaluation, while largely preserving performance on other benchmarks.
@@ 2.1 Characteristics of Human-likeness From Turing Tests
The Turing test is an imperfect proxy for human-likeness, but it provides two ingredients that are difficult to obtain otherwise: paired dialogues that are designed to be compared, and which side is human. We analyze the Turing test dialogue transcripts released by Jones and Bergen (2025).
@@
Running the LLM-judge over the dataset yields $2735$ total natural-language reason statements, many of which are redundant. We embed each reason statement using a sentence encoder Wang et al. (2020) and cluster them using a density-based clustering algorithm Campello et al. (2013). We find 53 representative clusters and extract their centroid. The 53 centroid statements still contained redundancies. Hence, we further instruct an LLM to summarize the 53 clusters into distinct Likert-style statements (prompt Appendix Figure 7). This yields a final inventory of 32 characteristics, presented in Appendix Table 5. Henceforth, we refer to these 32 traits as Human-Like 32 Questions or HL32Q.
@@ 3.1 Human-likeness Classifier
Given a dialogue, an LLM judge (HL32Q Judge) rates its agreement with each of the 32 statements only based on the witness responses on a 1–5 Likert scale (prompt Appendix Figure 10). Unlike the Turing test setting, this is not a pairwise comparison. Each dialogue is scored independently, producing a fixed-length feature vector $\mathbf{A}\in\mathbb{R}^{32}$.
@@ 3.2 Feature Reduction and Single Score of Human-likeness
Table 2: Finding characteristics aim to identify the differences between human–human and human–AI data and generate plausible reasons for these differences. The HL32Q judge aims to determine optimal weights for calculating a numerical human-likeness score. Accuracy on the filtered Turing test dataset from Jones and Bergen (2025) serves as a proxy for both tasks.
@@ 3.3 Evaluating on OOD Data
Finally, we test whether the HL16Q score generalizes beyond the Turing test data used to derive it. We evaluate on an out-of-distribution dataset consisting of 73 human–human and 73 human–AI dialogues from a separate cancer communication study Haut et al. (2025). This dataset differs in topic, style, and collection procedure (Data proxy in Appendix Table 6).
@@ 4 Inducing Human-likeness with Alignment
We frame this as a preference learning problem. For a given conversational prompt, we generate multiple candidate dialogues, score them with the HL16Q Judge, and construct ranked pairs where the more human-like dialogue is preferred. We use these pairs to align models with Direct Preference Optimization (DPO) Rafailov et al. (2023).
@@ 4.1 Persona Synthesis
To create diverse yet controlled conversational settings, we synthesize personas that serve as prompts for dialogue generation. We begin with 500 seed personas from the SynthLabs PERSONA dataset Castricato et al. (2025). These seeds are split into 450 training personas, 25 test personas, and 25 personas reserved for human evaluation. All augmentation is performed after this split to avoid data leakage.
@@ 4.3 Training
We fine-tune seven open-access models from multiple model families and generations, with parameter counts of 1B, 3B, 8B, 14B, 32B, 70B, and 72B, using DPO Rafailov et al. (2023). The models were simple instruction-tuned models, with no Mixture-of-Experts (MoE) and Chain-of-Thought (CoT) training, as these add more complexity in alignment training. Unlike standard alignment pipelines, we do not include an intermediate supervised fine-tuning (SFT) step, which typically helps the model learn the data format. All models reliably followed the required dialogue format with prompting alone, making SFT not essential.
We train all models for 10 epochs with DPO using $\beta=0.1$, AdamW optimization in 8-bit precision, a learning rate of $5\times 10^{-5}$, linear scheduling, and a 10% warmup. Training uses an effective batch size of 32 via gradient accumulation, a maximum sequence length of 1024, and no weight decay. We apply LoRA Hu et al. (2022) with rank 16, $\alpha=32$, and dropout 0.1. All models are trained with 4-bit quantization using HuggingFace Accelerate huggingface.co (2026) in data-parallel mode on 4 NVIDIA H100 GPUs.
@@ 4.4 Training Validation
Despite these improvements, mean HL16Q scores remain negative for most models after training. We attribute this to a domain mismatch between the alignment data and the original Turing test data, for which the bias $b$ was set.
@@ 5.1 Chatbot Arena
We adopt a Chatbot Arena–style Chiang et al. (2024) interface for pairwise comparison. Using the held-out set of 100 personas, participants interact with two chatbots displayed side by side in random order (Figure 5). For each trial, two models are randomly selected and both receive the same persona prompt. Participants are required to converse with each chatbot for at least two turns before making a decision from 5 choices: Certainly A, Likely A, Tie, Likely B, Certainly B.
@@ Results.
We report both win-rate and Elo scores, using a modified Elo system adapted from Chatbot Arena lmsys.org (2026) that supports partial wins. Because Elo is sensitive to comparison order, we report the mean Elo score over 500 random shuffles. Full details of the scoring procedure are provided in Appendix A.
@@ 5.2 Emotional Intelligence Benchmarks
To examine whether alignment for human-likeness degrades performance on other capabilities, we evaluate models on two widely used emotional intelligence benchmarks: EmoBench Sabour et al. (2024) and EQBench3 Paech (2023, 2025). We report results before and after alignment for all models, with full benchmark breakdowns provided in Appendix Table 9.
@@ 7 Limitations
A second limitation concerns the cost of computing HL16Q scores, which requires direct calls to a judge model. Although this cost is incurred only once during data construction, it is significantly higher than training a lightweight reward model, as in RLHF Ouyang et al. (2022), or using programmatic rewards in domains such as coding or mathematics Shao et al. (2024). While HAL enables alignment with richer and more nuanced rewards, this judging cost limits scalability at very large scale. In practice, the trade-off between reward expressiveness and computational cost must be considered when deciding whether HAL is suitable for a given alignment task. A promising direction for future work is to distill the HL16Q judge into smaller or ensemble reward models that retain performance while substantially reducing inference cost.
@@ Appendix B Additional Tables
Table 6: OOD dataset Data Proxy, showing structural comparison of Human–AI and Human–Human conversations in a medical setting. Here, the doctor serves as the investigator (I), who interacts with human patient actors and an AI patient as the witness (W). Full data cannot be released due to IRB protection. Although the goal in this study was not to differentiate human witnesses from AI, the pairwise data makes it suitable for our validation. More details about this data at Haut et al. (2025).